### v1

@@ Abstract
Numerical simulations of extrusion-based printing require tracking evolving material boundaries, a challenging task due to possible topological changes and mass conservation issues. Inaccurate conservation of mass can lead to a mismatch between the extruded and simulated shapes, and generally to unreliable predictions of the actual ink behavior. This work investigates the mass conservation properties of the conservative level-set method in extrusion-based 3D printing applications. We analyze the effects of the level set parameters on the accuracy of mass conservation using the cross-sectional area of the deposited strand. We compare the cross-sectional areas obtained in the simulation with the ideal areas obtained from a mass balance when the system reaches a steady-state condition. The numerical results indicate that reducing the reinitialization and the interface thickness parameters decreases the errors in the cross-sectional area obtained. However, the reductions in error tend to decline and could lead to excessive computational cost. Furthermore, we also found that the typical strong mesh requirements can be lessened by selecting an adequate interface thickness. Finally, we obtained the cross-sectional areas from simulations with different printing settings and found that they show good agreement with the simulated and experimental data published in previous work.
@@ 1 INTRODUCTION
Extrusion-based additive manufacturing (EB-AM) has gained widespread application due to its relative simplicity and material flexibility. This 3D printing process can accommodate various materials (e.g., food, ceramics, bioinks, or concrete) and geometric scales (from $\mu m$ to $cm$), offering versatility in high- and low-volume manufacturing applications Altıparmak et al. (2022). However, this versatility must be matched with a reliable and consistent process. This is particularly critical when working with materials where experimental iteration may be limited or expensive Naghieh and Chen (2021). To address these limitations, simulations can predict and optimize the behavior of the process, reducing the cost and variability of exclusive physical experiments. Consequently, numerical modeling and simulation of the process have become fundamental tools for understanding how machine settings and material properties affect a printed object Balta and Altınkaynak (2022).
A well-defined simulation should accurately represent the essential physical mechanisms of the printing process. This implies selecting the appropriate properties of the material, boundary conditions, and numerical methods. A commonly applied strategy for two-phase flow simulations is the one-fluid method Tryggvason (2011). This approach employs a single set of equations for the entire domain but requires the location of the phases to compute the necessary properties (e.g., density and viscosity) Tryggvason (2011). A consistent definition of the fluid interface is critical to enforce first principles, such as mass conservation, and handle discontinuities throughout the domain Mirjalili et al. (2017). For 3D printing simulations, the most prevalent interface-capturing methods in the literature are the volume of fluid (VOF) Spangenberg et al. (2021); Mollah et al. (2021), the coupled level set with VOF (CLSVOF) Comminal et al. (2018); Serdeczny et al. (2019) and the conservative level set (CLS) Balta and Altınkaynak (2022); Bakrani Balani et al. (2023). The VOF methods are based on sharp interface capturing, while the CLS considers a diffuse transition. Sharp capturing methods require a more complex numerical solution due to the jumps in the interface, but have good conservation properties. In contrast, the CLS approach is easier to implement, but can have issues with mass conservation Mirjalili et al. (2017).
This study is centered on the CLS method. The CLS approach was proposed to maintain the simplicity of the classical level set method while improving its conservation capabilities Olsson and Kreiss (2005). The authors proposed the use of a smooth level set function and a specific advection scheme to induce good conservation properties. However, a reinitialization step is necessary to maintain the thickness of the interface constant Olsson and Kreiss (2005). This reinitialization can affect the convergence of the solution Olsson et al. (2007) and requires careful treatment to avoid numerical accumulation of errors in the case of complex flows or long simulations Desjardins et al. (2008); McCaslin and Desjardins (2014); Chiodi and Desjardins (2017). In Chiu and Lin (2011), a more computationally attractive CLS strategy was proposed. The advection and reinitialization steps were combined into a single equation, reducing the computational demand in the original CLS. Although skipping the additional step (i.e, reinitialization) is desired, this method requires simultaneous adjustment of the parameter controlling the interface thickness $\epsilon$ and the amount of reinitialization induced $\gamma$.
In this work, we analyze the effects of the conservative level set parameters ($\epsilon,\gamma$) on the enforcement of mass conservation. This is quantified through the cross-sectional area of the printed strand. Different criteria have been proposed to select CLS parameters based on the local speed of the interface and the size of the mesh Jain (2022); Chiu and Lin (2011); Mirjalili et al. (2020). However, these criteria have been proposed on the basis of specific numerical discretization and meshing conditions. In 3D printing simulations, it is important to correctly capture the shape of the strand to properly characterize the roughness of the surface and the bonding between adjacent layers Comminal et al. (2018). Although there have been some works where the areas of a strand have been studied and validated against experiments Comminal et al. (2018); Bakrani Balani et al. (2023), this is the first time a numerical study of mass conservation has been conducted for 3D printing simulations. The remainder of this paper is organized as follows. In section II, we present all the relevant details for the modeling and simulation. Section III describes the methodology of the parametric analysis and the convergence test. The effects of the level set parameters are discussed in Section IV. Finally, the conclusions of this work are given in Section V.
@@ 2 MODEL DESCRIPTION
2.1 Fluid Flow Governing Equations
$\mathbf{\nabla}\cdot\mathbf{v}=0$ (1)
$\rho\left(\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\mathbf{\nabla})\mathbf{v}\right)=-\mathbf{\nabla}\cdot\left(-p\mathbf{I}+\mathbf{\sigma}\right)+\mathbf{F}_{o}$ (2)
where $\rho$ denotes density, $\mathbf{v}$ velocity, $p$ pressure, $\sigma$ the stress tensor and $\mathbf{F}_{o}$ the force contribution from gravity and (or) surface tension. A generalized stress tensor is defined as:
$\sigma=2\mu\mathbf{D}$ (3)
here $\mu$ is the viscosity and $\mathbf{D}$ is the strain-rate tensor. The viscosity of the ink depends on various factors, such as the properties of the material and the processing conditions, and is characterized using rheological testsWilms et al. (2021). A Newtonian constitutive model can be used when the main objective of the model is to predict the cross-sectional shape of the strand Serdeczny et al. (2018b). However, more complex models could be employed to predict the printing force Serdeczny et al. (2018b) or to include the effects of thermal conditions on the process Bakrani Balani et al. (2023).
2.2 Conservative Level Set Method
To represent the laminar two-phase (air-ink) flow that takes place during the deposition, a scalar field $\phi$ is introduced to model the interface between the two fluids as shown in Fig 1. For the level set method, $\phi$ varies between 0 (blue-ink) and 1 (red-air) and is advected by the velocity (computed with (1) and (2)) as follows:
$\frac{\partial{\phi}}{\partial t}+\mathbf{\nabla}\cdot(\mathbf{v}\phi)=F$ (5)
$F=\gamma\mathbf{\nabla}\cdot\left(\epsilon\nabla\phi-\phi(1-\phi)\frac{\nabla\phi}{|\nabla\phi|}\right)$ (6)
where $\epsilon$ is a parameter controlling the interface thickness and $\gamma$ is the reinitialization parameter Chiu and Lin (2011). The conservative level set method proposed in Olsson and Kreiss (2005) has a relatively simple structure and is especially suitable for simulations of incompressible flow, because Eq. 5 has the form of a conservative law. If ideal advection is considered, then $F=0$. For some specific initial conditions of $\phi$, the analytical solution of Eq. 5 is equivalent if $F=0$ or if $F$ is given by Eq. 6. However, as shown in Olsson et al. (2007), it is necessary to stabilize Eq. 5 because numerical perturbations are not damped and could create distorted shapes during the solution. The effect of perturbations on the solution can be studied by recasting the level-set evolution as the Burgers equation. In Olsson et al. (2007), it was found that the effect of numerical perturbations on the solution has a time scale $t\sim{\epsilon}/{\gamma}$. Based on this time scale, the criteria to select $\epsilon$} and $\gamma$} are normally supported by the assumption that if ${\epsilon}/{\gamma}<<1$ then the perturbation effects would be significantly faster than the advection dynamics. In this work, we use the single-step CLS method, which does not require a separate reinitialization step Chiu and Lin (2011). If Eq. 5 is solved using conservative numerical schemes and the level set parameters $\epsilon$ and $\gamma$ are properly tuned, then the mass of the fluids should be conserved.
There are some empirical recommendations on how to tune the level-set parameters. The thickness of the interface $\epsilon$ is normally selected around the maximum element size of the mesh Olsson et al. (2007); Chiu and Lin (2011). Regarding the reinitialization parameter, the recommendations in Olsson et al. (2007) are to use $\gamma=1$ for constant velocities (for a small $\epsilon$) or, based on the velocity gradient, to make $\gamma$ equal to the magnitude of the maximum velocity ($v_{max}$). Furthermore, for specific numerical methods and mesh definitions, $\phi$ is bounded for $\gamma>v_{max}$ and $\epsilon>0.5\Delta x$, where $x$ is the cell size of the mesh Jain (2022). Similarly, in Mirjalili et al. (2020), the authors found that, rather than using independent specific values, there should be a balance between $\epsilon$ and $\gamma$.
2.3 Boundary and Coupling Conditions
The boundary conditions of the model are the no-slip condition on the nozzle walls, symmetry around the $xz$ plane, an inlet flow rate on the top of the cylinder, the sliding wall at the bottom of the rectangular block, and open boundaries for the rest of the walls. From an operational perspective, two boundary conditions deserve more attention because they incorporate the manipulated variables of the process as follows:
$V_{pl}=-\int_{\partial\Omega_{inl}}\mathbf{v}\cdot\mathbf{n}\,dS,$ (7)
$\mathbf{v}_{\partial\Omega_{bn}}=\mathbf{v}_{bn}\quad\mathbf{v}_{bn}=(\bar{v}_{x},0,0)$ (8)
In Eq 7, the normal contributions of $\mathbf{v}$ over the boundary $\partial\Omega_{inl}$ are given by the flow rate $V_{pl}$, and in Eq 8, the velocity on the boundary $\partial\Omega_{bn}$ is set to $\bar{v}_{x}$, the relative nozzle speed. In terms of control, $\bar{v}_{x}$ can be directly adjusted using the G code, but $V_{pl}$ is implicitly related to the dynamics of the extrusion system.
$\rho=\rho_{1}+(\rho_{2}-\rho_{1})H(\phi)$ (9)
$\mu=\mu_{1}+(\mu_{2}-\mu_{1})H(\phi)$ (10)
where $H$ is a smooth Heaviside function Olsson et al. (2007); Chiu and Lin (2011).
2.4 Simulation Settings
[…]
((a))
((b))
Figure 1: (a) Simulation domain. (b) Volume fraction of fluid (blue-ink, red-air
@@ 2.4 Simulation Settings
The simulation is implemented in COMSOL Multiphysics® v6.2 COMSOL (2024), using the Laminar Flow and Level Set interfaces coupled with the Two-Phase Flow Level Set multiphysics. A stationary study step is defined to have a consistent initialization of the time-dependent solution, and the boundary conditions are ramped up to avoid numerical issues. The domain is defined according to the work in Serdeczny et al. (2018a). The extrusion nozzle is modeled as a cylinder with diameter $D$ and length $L$, and the distance between the tip of the nozzle and the bed is $\Delta Z$. The deposition domain is represented by a rectangular block, with cross-sectional dimensions $B_{w}\times B_{h}$ and length $B_{l}$. A three-dimensional representation of the nozzle considering the $xz$ symmetry of the domain is depicted in Fig 1. The effects of gravity are considered in the simulations and, for simplicity, the material is represented by a Newtonian constitutive model with viscosity $\mu$. All properties and dimensions used in the simulation are given in Table 1.
Table 1: Simulation Parameters

### v2

@@ Abstract
Accurate numerical simulation of material extrusion additive manufacturing requires reliable tracking of evolving material interfaces while preserving mass conservation. Inaccurate mass conservation can lead to significant discrepancies between simulated and deposited strand geometries, undermining the predictive capability of the model. In this work, we investigate the mass conservation performance of the conservative level-set (CLS) method in extrusion-based 3D printing simulations. A systematic parametric study is conducted to quantify the influence of the interface thickness and reinitialization parameters on mass conservation, using the steady-state cross-sectional area of deposited strands as a quantitative metric. Simulated cross-sections are compared against reference values obtained from analytical mass balance relations. The results show that reducing both the interface thickness and the reinitialization parameter improves mass conservation accuracy, although diminishing returns and increased computational cost are observed beyond certain thresholds. In addition, appropriate tuning of the interface thickness can relax mesh refinement requirements while maintaining acceptable accuracy. The proposed parameter selection strategy is validated across a range of printing conditions, materials, and nozzle geometries, including multilayer deposition of viscoplastic fluids. The simulations show reasonable agreement with experimentally validated data from the literature, confirming that careful CLS parameter tuning enables accurate and computationally efficient prediction of strand geometry in extrusion-based 3D printing.
@@ 1 INTRODUCTION
Extrusion-based 3D printing, also known as Material Extrusion Additive Manufacturing (MEX-AM), has become widely adopted due to its ability to process a diverse set of materials, including bioinks, ceramics, food, and concrete, across scales ranging from micrometers to centimeters (6_altiparmak_extrusion-based_2022). Despite this versatility, achieving a reliable and consistent process remains challenging, particularly when experimental iteration is costly or limited (7_naghieh_printabilitykey_2021). As a result, numerical simulation has been recognized as an effective means of investigating how material properties (e.g., viscosity, yield stress, and surface tension) and processing conditions (e.g., extrusion pressure, nozzle velocity, curing, and layer height) affect strand formation, surface quality, and interlayer bonding (3_balta_numerical_2022; 11_mollah_stability_2021).
Simulating MEX-AM requires multiphysics modeling of a two-phase flow, in which the extruded material interacts with the surrounding air and the fluid interface evolves over time. A consistent definition of the interface is therefore essential for enforcing mass conservation and managing discontinuities, making its numerical treatment a critical element of the solution strategy (5_Mirjalili2017_interface_capturing). Choosing the right approach for interface modeling is therefore a key step in ensuring accurate and reliable simulation results. Depending on how the computational mesh is used, interface-handling techniques can be broadly categorized into interface-capturing and interface-tracking methods, with the former adopting a Eulerian description on a fixed mesh and the latter relying on a Lagrangian or hybrid approach that follows the interface motion (elgeti_deforming_2016). In 3D printing simulations, the most popular solutions are based on interface-capturing methods, including the volume of fluid (VOF) (10_spangenberg_numerical_2021; 11_mollah_stability_2021), the coupled level-set with VOF (1_comminal_numerical_2018; 9_serdeczny_numerical_2019) and the conservative level-set (CLS) (3_balta_numerical_2022; 8_bakrani_balani_investigation_2023). The primary distinction between these methods lies in how they represent the interface. VOF methods employ a sharp interface, which ensures good mass conservation but requires more complex numerical treatment to handle discontinuities. In contrast, the CLS method uses a continuous, diffuse interface that is simpler to implement but requires careful parameter tuning to maintain mass conservation (5_Mirjalili2017_interface_capturing).
Among the interface-capturing methods, the conservative level-set (CLS) approach is particularly appealing for MEX-AM simulations because it combines a smooth interface representation with improved conservation properties. The CLS approach was originally proposed to retain the simplicity of the classical level-set method while improving its mass conservation capabilities (12_olsson_conservative_2005). This formulation employs a smooth level-set function along with a specific advection scheme to enhance conservation. However, a reinitialization step is required to maintain a constant interface thickness, which can affect convergence and, in complex flows or long simulations, lead to error accumulation (15_desjardins_accurate_2008; 14_mccaslin_localized_2014; 16_chiodi_reformulation_2017). To address this limitation, a more computationally efficient variant (18_chiu_conservative_2011) combines the advection and reinitialization steps into a single equation, reducing computational demand, although it requires careful selection of the interface thickness and the reinitialization parameters. While the CLS method offers distinct advantages, its predictive accuracy depends critically on parameter selection. Therefore, accurate selection of the CLS parameters is crucial in MEX-AM simulations, as it could directly influence predictions of strand cross-section, surface roughness, and interlayer adhesion.
Although previous studies have evaluated the influence of process settings on strand geometries via simulations validated with experiments (1_comminal_numerical_2018; 8_bakrani_balani_investigation_2023; 3_balta_numerical_2022), systematic analysis of parametric effects, especially regarding mass conservation, has been lacking. Given that mass conservation can directly affect the quality of predictions, understanding the CLS parameters’ influence is essential for accurate simulation. In this work, we investigate the influence of CLS parameters on mass conservation and strand formation, providing quantitative metrics to assess simulation accuracy. We employ the conservative level-set method to simulate multilayer MEX-AM, extending the formulation with a moving-mesh approach to accommodate vertical motion of the substrate. The simulation results are validated against experimental and simulation data for single and multilayer deposition with different materials and nozzle geometries.
The remainder of this article is organized as follows. In Section II, we present all the relevant details for the modeling and simulation. Section III describes the methodology of the parametric analysis and the single and multilayer study cases. In Section IV, we discuss the effects of the level-set parameters and evaluate the parameter selection for multiple case studies. Finally, the conclusions of this work are given in Section V.
@@ 2 MODEL DESCRIPTION
The model used in this study is based on a one-fluid formulation, in which the dynamics of the flow is governed by a single set of Navier-Stokes equations and the interface between phases is captured using the level-set method. Within this framework, the single-step CLS method tracks the evolving interface, with coupling achieved through the advection velocity and averaging of phase properties. In-plane motion is imposed through a boundary condition; however, multilayer printing introduces relative vertical motion, which requires updating one of the domain boundaries. To accommodate this, the formulation is extended with a moving-mesh approach (arbitrary Lagrangian–Eulerian), allowing the material frame to be applied locally while maintaining accurate interface tracking and conservation.
Figure 1: a) Boundary conditions. b) Single-layer simulation domain. c) Multilayer simulation domain
2.1 Single-layer simulations
@@
The domain for single-layer simulations was defined according to the work in 2_serdeczny_experimental_2018 and a three-dimensional representation of the domain considering the $xz$ symmetry is depicted in Fig 1b. The nozzle was modeled as a cylinder with diameter $D$ and length $L$, and the distance between the tip of the nozzle and the bed is $\Delta Z$. The deposition domain was represented by a rectangular block, with cross-sectional dimensions $B_{w}\times B_{h}$ and length $B_{l}$. The effects of gravity were considered in the simulations and all properties and dimensions used in the single-layer simulations are given in Table 1.
Table 1: Single layer - Simulation Parameters
@@
2.2 Multilayer simulations
The multilayer simulations followed the printing settings reported by el_abbaoui_3d_2024, 10_spangenberg_numerical_2021 and co_comminal_modelling_2020 for the deposition of concrete using a cylindrical nozzle, and by Mollah2025RapidCuringCFD for the deposition of metakaolin using a rectangular nozzle.The nozzle was placed in the center of the deposition domain, and the length $B_{l}$ was increased to accommodate the horizontal motion in both directions. The simulation domain for the rectangular nozzle is shown in Figure 1c, and the domain for the cylindrical nozzle followed the same configuration. The smoothing of the moving mesh is based on the Yeoh method with a stiffening factor $C_{2}=100$. The boundaries defining the nozzle were fixed, and the substrate wall had a prescribed horizontal speed (printing speed) and a normal mesh velocity (vertical speed), as shown in Figure 2.
Figure 2: Printing and vertical speeds a) Concrete b) Metakaolin
The cylindrical nozzle has a diameter $D=25$ mmmm and the rectangular nozzle has a cross section of $20\times 4$ $mm^{2}$. All the properties and dimensions used for the multilayer simulations are given in Table 2.
Table 2: Multilayer - Simulation Parameters
Parameter Nomenclature Values (Concrete) Values (Metakaolin) Units
$L$ Nozzle Length 150 67 mm
$\Delta Z$ Nozzle to bed distance 12.5 3.6 mm
$B_{w}$ Block width 80 40 mm
$B_{h}$ Block height 50 14.4 mm
$B_{l}$ Block Length 630 150 mm
$V_{pl}$ Inlet flow 8240 180 mm3/s
$\rho$ Density of the material 2.1 1.63 g/cm3
$\mu_{p}$ Plastic Viscosity 7.5 0.4746 Pa·s
$\tau_{y}$ Yield stress 630 41.69 Pa
$m_{p}$ Bingham regularization 100 100 s
2.3 Fluid flow equations
$\mathbf{\nabla}\cdot\mathbf{v}=0$ (1)
$\rho\left(\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\mathbf{\nabla})\mathbf{v}\right)=\mathbf{\nabla}\cdot\left(-p\mathbf{I}+\mathbf{\sigma}\right)+{\mathbf{F}_{g}+\mathbf{F}_{s}}$ (2)
where $\rho$ denotes density, $\mathbf{v}$ velocity, $p$ pressure, $\sigma$ the deviatoric stress tensor, $\mathbf{F}_{g}$ the body force term from gravity and $\mathbf{F}_{s}$ the force due to surface tension. A generalized stress tensor is defined as:
$\sigma=2\mu\mathbf{D}$ (3)
here $\mu$ is the viscosity and $\mathbf{D}$ is the strain-rate tensor. The viscosity of the ink is governed by its composition and microstructure, together with processing conditions such as shear rate, temperature, and printing parameters, and is characterized using rheological tests (wilms_formulation_2021). As reported in 20_Serdeczny2018_shear_thinning, a Newtonian model can be employed when the primary objective is to predict the cross-sectional shape of the deposited strand. In contrast, more complex models can be adopted to capture more complex effects, such as elasticity and yield stress, which improve the prediction of printing forces (serdeczny_viscoelastic_2022). In this study, a Newtonian constitutive model with viscosity $\mu=\mu_{s}$ is used for the single-layer simulations, and the printing materials for the multilayer simulations are modeled using the Bingham-Papanastasiou constitutive model:
$\mu_{m}=\mu_{p}+\frac{\tau_{y}}{\dot{\gamma}}(1-\exp(-m_{p}\dot{\gamma}))$ (5)
In this case $\mu=\mu_{m}$ and the model is parameterized by the plastic viscosity $\mu_{p}$, the yield stress $\tau_{y}$, the shear rate $\dot{\gamma}$, and the regularization parameter $m_{p}$.
2.4 Conservative level-set method
The conservative level-set method proposed in 12_olsson_conservative_2005 introduces a continuous scalar field $\phi\in[0,1]$ that defines the interface between two fluids as the contour $\phi=0.5$. The evolution of the interface between the two phases requires the advection of $\phi$ based on the velocity $\mathbf{v}$ and is given by the following conservative form:
$\frac{\partial{\phi}}{\partial t}+\mathbf{\nabla}\cdot(\mathbf{v}\phi)=0$ (6)
This expression has a simple structure and is especially suitable for simulations of incompressible flow. Nevertheless, as shown in 13_olsson_conservative_2007, it is necessary to stabilize Eq. (6) because numerical perturbations are not damped and create distorted shapes during the solution. 13_olsson_conservative_2007 formulated a stabilized advection based on an additional term that depends on a reinitialization parameter $\gamma$} and a parameter controlling the thickness of the interface $\epsilon$,
$\frac{\partial{\phi}}{\partial t}+\mathbf{\nabla}\cdot(\mathbf{v}\phi)=\gamma\mathbf{\nabla}\cdot\left(\epsilon(\nabla\phi\cdot\hat{n})\hat{n}-\phi(1-\phi)\hat{n}\right),\qquad\text{where}\quad\hat{n}=\frac{\nabla\phi}{|\nabla\phi|}.$ (7)
However, the solution of this equation is usually divided into advection and stabilization steps for numerical reasons. In this work, we use the single-step CLS method (18_chiu_conservative_2011), which does not require a separate stabilization step. The advection of the level-set function is based on the following expression:
$\frac{\partial{\phi}}{\partial t}+\mathbf{\nabla}\cdot(\mathbf{v}\phi)=\gamma\mathbf{\nabla}\cdot\left(\epsilon\nabla\phi-\phi(1-\phi)\frac{\nabla\phi}{|\nabla\phi|}\right)$ (8)
The only difference between the stabilized advection (Eq. (7)) and the single-step level-set evolution (Eq. (8)) is the projection of the gradient $\nabla\phi$ along the normal direction $\hat{n}$. Since the single-step CLS method requires the solution of a single PDE, its solution is computationally more efficient than the original conservative level-set method. Furthermore, if Eq. (8) is solved using conservative numerical schemes and the level-set parameters $\epsilon$ and $\gamma$ are properly tuned, then the mass of the fluids should be conserved.
A variety of empirical and analytical guidelines have been proposed for tuning the level-set parameters. 13_olsson_conservative_2007 showed that the effect of numerical perturbations on the solution evolves on a time scale of $t\sim\epsilon/\gamma$. Based on this timescale, it is assumed that if $\epsilon/\gamma\ll 1$, perturbation effects are significantly faster than the advection dynamics. The interface thickness $\epsilon$ is typically chosen to be on the order of the maximum element size of the mesh (13_olsson_conservative_2007; 18_chiu_conservative_2011). Regarding the reinitialization parameter, 13_olsson_conservative_2007 recommended using $\gamma=1$ for constant velocities (when $\epsilon$ is small), or setting $\gamma$ according to the velocity gradients at the interface. 18_chiu_conservative_2011 selected $\gamma$ to match the maximum velocity in the domain. Other studies have proposed criteria for selecting $\epsilon$ and $\gamma$ based on the stability of the numerical solution and the boundedness of the level-set field $\phi$ (19_mirjalili_conservative_2020; 172_jain_conservative_2020). These parameters have also been shown to constrain the allowable time-step size, potentially increasing the computational cost (17_jain_accurate_2022). However, these analyses are typically based on uniform meshes and central finite difference schemes, which limits their direct applicability to other discretization methods.
2.5 Boundary and coupling conditions
The boundary conditions of the model are the no-slip condition on the nozzle walls, symmetry around the $xz$ plane, an inlet flow rate on the top of the cylinder, the sliding wall at the bottom of the rectangular block, and open boundaries for the rest of the walls in the deposition domain. The domain and the boundaries considered in a general 3D printing simulation are represented in Figure 1. From an operational perspective, two boundary conditions deserve more attention because they incorporate the manipulated variables of the process as follows:
$V_{pl}=-\int_{\partial\Omega_{inl}}\mathbf{v}\cdot\mathbf{n}\,dS,$ (9)
$\mathbf{v}_{\partial\Omega_{bn}}=\mathbf{v}_{bn},\quad\mathbf{v}_{bn}=(\bar{v}_{x},0,0)$ (10)
In Eq 9, the normal contributions of $\mathbf{v}$ over the boundary $\partial\Omega_{inl}$ (nozzle inlet) are given by the inlet flow rate $V_{pl}$, and in Eq 10, the velocity on the boundary $\partial\Omega_{bn}$ (substrate) is set to $\bar{v}_{x}$, the relative printing speed. In terms of control, $\bar{v}_{x}$ can be directly adjusted using the G code, but $V_{pl}$ is implicitly related to the dynamics of the extrusion system.
$\rho=\rho_{1}+(\rho_{2}-\rho_{1})H(\phi)$ (11)
$\mu=\mu_{1}+(\mu_{2}-\mu_{1})H(\phi)$ (12)
where $H$ is a smooth Heaviside function (13_olsson_conservative_2007; 18_chiu_conservative_2011).
2.6 Moving mesh
Two-phase flow simulations based on the one-fluid formulation require a multiphysics approach to capture the interface between the fluids. The horizontal motion of the strand can be imposed using a sliding wall boundary condition on the substrate. However, a multilayer printing simulation requires a relative vertical displacement of the bed (or the nozzle), which cannot be incorporated by using a kinematic condition at the boundary. Since a purely Eulerian formulation cannot accommodate moving boundaries, a reformulation of the equations is needed. Therefore, we define a moving mesh to handle the change in the domain when the system has the aforementioned vertical motion. This feature is based on an arbitrary Lagrangian-Eulerian (ALE) method where the equations are expressed in a reference frame that follows the motion of the computational mesh. The mesh coordinates are dynamically updated to maintain mesh quality, with the mapping between the mesh and spatial coordinates adjusted as necessary (comsol).